# Multi-dimensional Evaluation of Empathetic Dialogue Responses

### v2

@@ Abstract
Empathy is critical for effective and satisfactory conversational communication. Prior efforts to measure conversational empathy mostly focus on expressed communicative intents—that is, the way empathy is expressed. Yet, these works ignore the fact that conversation is also a collaboration involving both speakers and listeners. In contrast, we propose a multi-dimensional empathy evaluation framework to measure both expressed intents from the speaker’s perspective and perceived empathy from the listener’s perspective. We apply our proposed framework to analyze our internal customer-service dialogue. We find the two dimensions (expressed intent types and perceived empathy) are inter-connected, and perceived empathy has a high correlation with dialogue satisfaction levels.
@@
Zhichao Xu1thanks:  Work done during ZX’s internship at Google Research  Jiepu Jiang2 1Kahlert School of Computing, University of Utah 2Google Research zhichao.xu@utah.edu
@@ 1 Introduction
Empathy as a complex socio-emotional phenomenon plays a crucial role in forming and maintaining social interactions (Omdahl, 2014). Prior works in psychology and social science have highlighted empathy’s importance in human communications (Omdahl, 2014; Goodwin and Heritage, 1990; Goodwin et al., 2012; Suchman et al., 1997, inter alia). As speech and text-based dialogues remain prevalent in human communication, developing computational approaches to analyze and measure empathy expressed in dialogues is gaining traction in the CL community (Rashkin et al., 2019; Liu et al., 2021; Sharma et al., 2020, inter alia).
@@
Prior efforts propose empathy frameworks grounded in psychology and social science theories. These frameworks are then adapted for specific applications or dataset domains. Correspondingly, researchers propose data-driven approaches to analyze and measure empathy. Upon reviewing existing works by the CL community, we found that they mostly adopt an individualistic conceptualization of empathy (Rashkin et al., 2019; Sharma et al., 2020; Welivita and Pu, 2020; Welivita et al., 2021) which position empathy as a primarily individual mental process. These works mainly focus on modeling expressed empathy in dialogues. However, recent psychology works (Concannon and Tomalin, 2023; Zaki et al., 2008) argue that empathy is more fundamentally relational: instead of a static personality trait, empathy is a dynamic process, aiming to serve interpersonal communicative purposes. Van Dijke et al. (2020) emphasize empathy as a collaborative practice, involves participation by both parties and is shaped in social interactions. Thus, studying perceived empathy in a collaborative, dynamic dialogue setting is of importance.
@@
In this study, we propose a novel multi-dimensional evaluation framework to assess perceived empathy in multi-round dialogues, measuring expressed empathetic intents alongside perceived empathy. Notably, our framework includes four perceived empathy dimensions: (1) perceived engagement, (2) perceived understand, (3) perceived sympathy and (4) perceived helpfulness. To validate this framework’s practicality, we annotate an internal dataset of 2,000 customer service dialogue sessions. and carry out a fine-grained analysis. Our findings show a strong correlation between perceived empathy and overall dialogue satisfaction ratings, affirming the framework’s efficacy in capturing empathetic traits.
@@
The proposed empathy evaluation framework still requires subjective assessments from human experts, which is expensive and non-trivial to collect. To scale up evaluation without excessive reliance on annotated data, we investigate two model-based evaluation approaches: prompting LLMs and training language model-based classifiers. We show that prompting methods with popular LLMs such as GPT-4 and Flan family models still perform poorly, indicating the challenging nature of measuring conversational empathy. In contrast, traditional supervised fine-tuning yields better accuracy with labeled data. Notably, instruction-finetuned classifiers based on Flan-T5 family models (Chung et al., 2022) achieve the best performance. We conduct a comprehensive ablation study of this instruction finetuning method to provide insights into its strong performance.
To summarize, our contributions are three-fold:
@@ 2 Related Work
Definition of empathy in literature.  Empathy has been an influential concept in ethics and moral theory (Hoffman, 1996; Slote, 2007; Smith, 1959), social/developmental psychology (Hoffman, 1996; Batson et al., 2015) and other domains. Contributions from various domains result in different definitions, functions and proposed empathy components, posing a challenge to studying empathy (Coplan, 2011). Prior works (Elliott et al., 2011, 2018; Althoff et al., 2016; Sharma et al., 2020, inter alia) mainly consider empathy as an individual’s mental process. Recent psychological studies highlight empathy as a collaborative practice involving participation by both parties (Van Dijke et al., 2020; Yalçın and DiPaola, 2020, inter alia). Within this work, we adopt this interactional definition of empathy, as our aim is to measure and evaluate empathy in social scenarios such as real-time human-human conversation.
@@ 3 Empathy Measurement Framework
Empathy as a collaborative practice.  Empathy is critical in human communication for social interactions (Omdahl, 2014), varying in definition across psychology and social science literature (Hall and Schwartz, 2019; Yalçın and DiPaola, 2020, inter alia). Elliott et al. (2018) distinguishes affective and cognitive aspects of empathy. The affective aspect relates to the emotional stimulation in reaction to the experiences and feelings expressed by others, while the cognitive aspect is a more deliberate process of understanding and interpreting the experiences and feelings of others and communicating such understandings. However, this interpretation overlooks empathy’s role in interactive dialogues where multiple parties engage. Recent psychology works empathize empathy as a collaborative, interactive process (Van Dijke et al., 2020; Yalçın and DiPaola, 2020; Concannon and Tomalin, 2023, inter alia), urging studies across diverse social scenarios. Motivated by these insights, our study adopts this interactional definition of empathy—empathy refers to an individual’s ability to perceive and understand the other’s emotional states and respond correspondingly.
@@
Rashkin et al. (2019) None Uni-dimensional
Sharma et al. (2020) 3 intents None
Welivita et al. (2021) 8 intents None
Liu et al. (2021) 7 intents None
Our framework 16 intents 4 dimensions
@@
  • Perceived Empathy: we also assess empathy from the listener’s perspective, predicting whether an utterance is perceived as empathetic. This assessment breaks perceived empathy into four fine-grained aspects:
    • Engagement: measures the degree to which the listener perceives the speaker as involved in the conversation. Low engagement can lead to the listener feeling indifferent and lack of empathy.
    • Understanding: measures how well the listener feels the speaker understands their situation, feelings or problems. A lack of understanding can diminish perceived empathy.
    • Sympathy: captures the listener’s perception of the speaker’s ability to empathize and react appropriately. A lack of sympathy can weaken perceived empathy.
    • Helpfulness: determines whether the listener finds the speaker’s communication to be helpful in addressing the conversation’s core issues. While not directly related to empathy, in task-oriented dialogues, where addressing problems is the objective, helpfulness can affect perceived empathy.
@@ 4 Empathy Measurement Models
Measuring conversational empathy is an intrinsically challenging task because: (1) dialogues vary in themes, types, and contexts; (2) empathy is often communicated implicitly (Suchman et al., 1997; Li et al., 2023; Concannon and Tomalin, 2023), posing challenges for state-of-the-art language models Ruis et al. (2022); (3) empathy expression styles and intents are often domain specific; and (4) the nuance between expressed and perceived empathy is often subtle and depends on the two parties’ social roles (Suchman et al., 1997; Deppermann, 2011; Kupetz, 2014, inter alia). For example, asking clarification questions may be perceived as empathetic in a dialogue between a therapist and a patient, but can be neutral in task-oriented dialogues such as customer service dialogues.
@@ 5.1 Datasets
ESConv (Liu et al., 2021) is collected with crowdworkers chatting in help-seeker and supporter roles, to simulate the scenario of synchronous peer-to-peer mental counseling and support. The dataset consists of 1.3k multi-round dialogues and is annotated with 3 mental support stages and 7 fine-grained mental support strategies. The dataset was originally designed for multi-class classification for the 7 communicative strategies (intents), and we adapt it to multi-label classification setting.
Empeval is our internal dataset (due to privacy concerns, we cannot release the annotated dataset, but we try to present a holistic view of the dataset to assist reproducbility). This dataset includes 1,833 anonymized conversations sampled from a company’s customer support logs, where the conversations are between customers and human representatives. All personal identifiers have been removed. On average, each conversation has 41.4 utterances, and each utterance has 7.4 tokens, as measured by T5 models’ WordPiece tokenizer. When we built the dataset, we randomly sampled 2,000 conversations from the log and used an internal human annotation platform to judge the conversations. The annotation questions are provided in Tables 10 and 11. Each conversation were annotated by two different raters recruited from an internal annotation platform. 167 non-English conversations were discarded. On average, the two raters have achieved 90.23% agreement rate and 0.615 Cohen’s kappa on the judgments—indicating a moderate rating consistency. For expressed communicative intents, we mark the intent as True when both annotators mark it as True. For the four perceived empathy dimensions, we convert the numerical annotations to binary labels by marking as True when both annotators give a score higher than 4.
@@ 5.2 Compared Methods
Sharma et al. (2020) use a bi-encoder model structure to jointly predict listener’s communicative intent and extract rationales. The model is initialized with RoBERTa-Base (Liu et al., 2019) with continued in-domain pretraining as discussed by (Gururangan et al., 2020). The method is designed for single-round conversations thus we also report the single-round performances on our two multi-round dialogue datasets ESConv and Empeval.
@@
5.3 Loss Functions
We notice longtail distributions of labels on all three datasets, due to the nature of human-human dialogues and categorization of communicative empathy. In addition to the standard cross-entropy loss function, we additionally include two other loss functions: Focal loss (Lin et al., 2017) and LDAM loss Cao et al. (2019) to tackle class imbalance problem. The main performance reported is still based on standard cross-entropy loss and we use these two additional loss functions as an ablation study (detailed in Section 6.2).
@@ 7 Conclusion and Future Work
This proposed evaluation framework requires subjective assessments from trained annotators, which can be expensive and non-trivial to collect. To explore automated empathy measurement models, we rigorously compare various methods against human judgments. Our instruction finetuning technique achieves the highest classification performance, measured by F-score and accuracy compared to human-based judgments. While our discussion primarily centers on measuring empathy perception in human-human conversations, an important future direction is extending this analysis to human-machine interactions.
@@
This paper studies the empathy evaluation problem in the context of human-human communication. A natural extension is to apply our framework to dialogue applications, and the findings in this paper may be subject to change in human-AI dialogues. Due to limited bandwidth, we only experiment with one proprietary model—GPT-4-0613. Including more proprietary models such as Gemini Team et al. (2023) and Claude will make the results more comprehensive.

### v3

@@ Abstract
Empathy is critical for effective and satisfactory conversational communication. Prior efforts to measure conversational empathy mostly focus on expressed communicative intents—that is, the way empathy is expressed. Yet, these works ignore the fact that conversation is also a collaboration involving both speakers and listeners. In contrast, we propose a multi-dimensional empathy evaluation framework to measure both expressed intents from the speaker’s perspective and perceived empathy from the listener’s perspective. We apply our analytical framework to examine internal customer-service dialogues. We find the two dimensions (expressed intent types and perceived empathy) are inter-connected, while perceived empathy has high correlations with dialogue satisfaction levels.
@@
Zhichao Xu1thanks:   Work done during ZX’s internship at Google Research.  Jiepu Jiang2 1 Kahlert School of Computing, University of Utah 2 Google Research zhichao.xu@utah.edu, jiepujiang@google.com
@@ 1 Introduction
Empathy as a complex socio-emotional phenomenon plays a crucial role in forming and maintaining social interactions (Omdahl, 2014). Prior works in psychology and social science have highlighted the importance of empathy in human communications (Omdahl, 2014; Goodwin and Heritage, 1990; Goodwin et al., 2012; Suchman et al., 1997, inter alia). As speech and text-based dialogues remain prevalent in human communication, many previous studies had developed computational approaches to analyze and measure empathy expressed in dialogues (Rashkin et al., 2019; Liu et al., 2021; Sharma et al., 2020, inter alia).
@@
Prior efforts had proposed empathy frameworks grounded in psychology and social science theories. These frameworks were then adapted for specific applications or dataset domains. Correspondingly, researchers proposed data-driven approaches to analyze and measure empathy. Upon reviewing existing works, we found that they mostly adopted an individualistic conceptualization of empathy (Rashkin et al., 2019; Sharma et al., 2020; Welivita and Pu, 2020; Welivita et al., 2021) which positions empathy as a primarily individual mental process. These works mainly focused on modeling expressed empathy in dialogues. However, recent psychology research (Concannon and Tomalin, 2023; Zaki et al., 2008) argued that empathy is more fundamentally relational: instead of a static personality trait, empathy is a dynamic process, aiming to serve interpersonal communicative purposes. Van Dijke et al. (2020) emphasized empathy as a collaborative practice, involving participation by both parties and is shaped in social interactions. Thus, it is critical to further examine perceived empathy in a collaborative, dynamic dialogue setting.
@@
In this study, we propose a novel conceptual framework to evaluate empathy in multi-round dialogues. Specifically, this framework measures both expressed and perceived empathy intents. Our proposed multi-dimensional evaluation framework includes four perceived empathy dimensions: (1) perceived engagement, (2) perceived understand, (3) perceived sympathy and (4) perceived helpfulness. To validate this framework’s practicality, we annotated an internal dataset of 2,000 customer service dialogue sessions. and performed a fine-grained analysis. Our findings show a strong correlation between perceived empathy and overall dialogue satisfaction ratings, affirming the framework’s efficacy in capturing empathetic traits.
@@
The proposed empathy evaluation framework still requires subjective assessments from human experts, which are expensive and non-trivial to collect. To scale up evaluation without excessive reliance on annotated data, we investigate two model-based evaluation approaches: prompting LLMs and training language model-based classifiers. We show that prompting methods with popular LLMs such as GPT-4 and Flan family models perform poorly on our datasets, indicating the challenging nature of measuring conversational empathy. In contrast, traditional supervised fine-tuning yields better accuracy with labeled data. Notably, instruction-finetuned classifiers based on Flan-T5 family models (Chung et al., 2022) achieve the best performance. We also conduct a comprehensive ablation study of this method to provide insights into its strong performance. To summarize, our contributions are three-fold:
@@ 2 Related Work
Definition of empathy in literature.  Empathy has been an influential concept in ethics and moral theory (Hoffman, 1996; Slote, 2007; Smith, 1959), social/developmental psychology (Hoffman, 1996; Batson et al., 2015) and other domains. Contributions from various domains result in different definitions, functions and proposed empathy components, posing a challenge to studying empathy (Coplan, 2011). Prior works (Elliott et al., 2011, 2018; Althoff et al., 2016; Sharma et al., 2020, inter alia) mainly consider empathy as an individual’s mental process. Recent psychological studies highlight empathy as a collaborative practice involving participation by both parties (Van Dijke et al., 2020; Yalçın and DiPaola, 2020; Concannon and Tomalin, 2023, inter alia). Within this work, we adopt this interactional definition of empathy, as our aim is to measure and evaluate empathy in social scenarios such as real-time human-human conversation.
@@ 3 Empathy Measurement Framework
Empathy as a collaborative practice.  Empathy is critical in human communication for social interactions (Omdahl, 2014), varying in definition across psychology and social science literature (Hall and Schwartz, 2019; Yalçın and DiPaola, 2020, inter alia). Elliott et al. (2018) distinguishes affective and cognitive aspects of empathy. The affective aspect relates to the emotional stimulation in reaction to the experiences and feelings expressed by others, while the cognitive aspect is a more deliberate process of understanding and interpreting the experiences and feelings of others and communicating such understandings. However, this interpretation overlooks empathy’s role in interactive dialogues where multiple parties engage. Recent psychology works empathize empathy as a collaborative, interactive process (Van Dijke et al., 2020; Concannon and Tomalin, 2023, inter alia), urging studies across diverse social scenarios. Motivated by these insights, our study adopts this interactional definition of empathy—empathy refers to an individual’s ability to perceive and understand the other’s emotional states and respond correspondingly.
  • Perceived Empathy: We adapt the definition in recent psychology literature (Concannon and Tomalin, 2023; Yalçın, 2019; Yalçın and DiPaola, 2020, inter alia) for task-oriented dialogues. We propose to assess empathy from the listener’s perspective, predicting whether an utterance is perceived as empathetic. This assessment breaks perceived empathy into four fine-grained aspects:
    • Engagement measures the degree to which the listener perceives the speaker as being involved in the conversation. Low engagement in the conversation can lead to the listener feeling indifferent and lacking empathy.
    • Understanding measures how well the listener feels the speaker understands their situation, feelings or problems. A lack of understanding can diminish perceived empathy.
    • Sympathy captures the listener’s perception of the speaker’s ability to empathize and react appropriately. A lack of sympathy can weaken perceived empathy.
    • Helpfulness determines whether the listener finds the speaker’s communication to be helpful in addressing the conversation’s core issues. While not directly related to empathy, in task-oriented dialogues, where addressing problems is the objective, helpfulness can affect perceived empathy.
@@
Rashkin et al. (2019) ✓(Uni-dimensional)
Sharma et al. (2020) ✓(3 intents)
Welivita et al. (2021) ✓(8 intents)
Liu et al. (2021) ✓(7 intents)
Our framework ✓(16 intents) ✓(4 dimensions)
@@ 4 Empathy Measurement Models
Measuring conversational empathy is an intrinsically challenging task because: (1) empathy is often communicated implicitly (Suchman et al., 1997; Li et al., 2023; Concannon and Tomalin, 2023), posing challenges for state-of-the-art language models Ruis et al. (2022); (2) empathy expression styles and intents are often domain specific; and (3) the nuance between expressed and perceived empathy is often subtle and depends on the two parties’ social roles (Suchman et al., 1997; Deppermann, 2011; Kupetz, 2014, inter alia). For example, asking clarification questions may be perceived as empathetic in a dialogue between a therapist and a patient, but can be neutral in task-oriented dialogues such as customer service dialogues.
@@ 5.1 Datasets
Emotion Support Conversation (ESConv) (Liu et al., 2021) is collected with crowdworkers chatting in help-seeker and supporter roles, to simulate the scenario of synchronous peer-to-peer mental counseling and support. The dataset consists of 1.3k multi-round dialogues and is annotated with 3 mental support stages and 7 fine-grained mental support strategies. The dataset was originally designed for multi-class classification for the 7 communicative strategies (intents), and we adapt it to multi-label classification setting.
Empathy Evaluation (Empeval) is our internal dataset (due to privacy concerns, we cannot release the annotated dataset, but we try to present a holistic view to assist reproducibility). This dataset includes 1,833 anonymized conversations sampled from a company’s customer support logs, where the conversations are between customers and human representatives. All personal identifiers have been removed. On average, each conversation has 41.4 utterances, and each utterance has 7.4 tokens measured by T5 models’ WordPiece tokenizer. When we built the dataset, we randomly sampled 2,000 conversations from the log and used an internal human annotation platform to judge the conversations. The annotation questions are provided in Tables 10 and 11. Each conversation were annotated by two different raters recruited from an internal annotation platform. 167 non-English conversations were discarded. On average, the two raters have achieved 90.23% agreement rate and 0.615 Cohen’s kappa on the judgments—indicating a moderate rating consistency. For expressed communicative intents, we mark the intent as True when both annotators mark it as True. For the four perceived empathy dimensions, we convert the numerical annotations to binary labels by marking as True when both annotators give a score higher than 4.
@@ 5.2 Compared Methods
Sharma et al. (2020) use a bi-encoder model initialized with RoBERTa-Base (Liu et al., 2019), to jointly predict listener’s communicative intent and extract rationales. The method is designed for single-round conversations thus we also report the single-round performances on our two multi-round dialogue datasets ESConv and Empeval.
@@
5.3 Evaluation Metrics
@@ 5.4 Implementation Details
Loss functions.  We notice longtail distributions of labels on all three datasets, due to the nature of human-human dialogues and categorization of communicative empathy. In addition to the standard cross-entropy loss function, we additionally include two other loss functions: Focal loss (Lin et al., 2017) and LDAM loss Cao et al. (2019) to tackle class imbalance problem. The main performance reported is still based on standard cross-entropy loss and we use these two additional loss functions as an ablation study (detailed in Section 6.2).
@@ 7 Conclusion and Future Work
This proposed conceptual evaluation framework requires subjective assessments from trained annotators, which can be expensive and non-trivial to collect. To explore automated empathy measurement models, we rigorously compare various methods against human judgments. Our instruction finetuning technique achieves the highest classification performance, measured by F-score and accuracy compared to human-based judgments. While our discussion primarily centers on measuring empathy perception in human-human conversations, an important future direction is extending this analysis to human-machine interactions.
@@
This paper studies the empathy evaluation problem in the context of human-human communication. A natural extension is to apply our framework to dialogue applications, and the findings in this paper may be subject to change in human-AI dialogues. Due to limited bandwidth, we only experiment with one proprietary model—GPT-4-0613. Including more latest proprietary models such as GPT-4o, Gemini Gemini et al. (2023), and Claude will make the results more comprehensive.